# PharmacoMatch: Efficient 3D Pharmacophore Screening through Neural Subgraph Matching

### v1

@@ Abstract
The increasing size of screening libraries poses a significant challenge for the development of virtual screening methods for drug discovery, necessitating a re-evaluation of traditional approaches in the era of big data. Although 3D pharmacophore screening remains a prevalent technique, its application to very large datasets is limited by the computational cost associated with matching query pharmacophores to database ligands. In this study, we introduce PharmacoMatch, a novel contrastive learning approach based on neural subgraph matching. Our method reinterprets pharmacophore screening as an approximate subgraph matching problem and enables efficient querying of conformational databases by encoding query-target relationships in the embedding space. We conduct comprehensive evaluations of the learned representations and benchmark our method on virtual screening datasets in a zero-shot setting. Our findings demonstrate significantly shorter runtimes for pharmacophore matching, offering a promising speed-up for screening very large datasets.
Keywords Contrastive representation learning  $\cdot$ Order embedding space  $\cdot$ Virtual screening  $\cdot$ Pharmacophore modeling
@@ 1 Introduction
A challenging task in the early stages of drug discovery campaigns is the identification of hit molecules that effectively bind to a protein target of interest. Due to the vastness of the chemical space, estimated to encompass more than $10^{60}$ small organic molecules (Virshup et al. 2013), identifying molecules with desirable drug-like properties is often compared to finding a needle in a haystack. Virtual screening methods have therefore become an essential component of the computer-aided drug discovery toolkit, aiding medicinal chemists in filtering molecular databases to efficiently explore the search space for potential hit compounds (Sliwoski et al. 2014).
A pharmacophore represents non-bonding interactions of chemical features that are essential for binding to a specific protein target (Wermuth et al. 1998). A pharmacophore query can, for example, be generated from the interaction profile of a ligand-receptor complex and used to identify potential hit compounds from databases by searching for molecules with similar pharmacophoric patterns (Wolber and Langer 2005). The process involves a positional alignment of the pharmacophore model with the three-dimensional conformations of molecules in the database, which are ranked based on their agreement with the pharmacophore query (Wolber, Dornhofer, and Langer 2006). Since pharmacophore screening focuses on abstract interaction patterns, rather than specific molecular structures, it allows for the identification of structurally diverse hit compounds (Seidel et al. 2017).
@@
Virtual screening of make-on-demand libraries like Enamine REAL (Shivanyuk et al. 2007) is of growing interest because these libraries contain compounds that can be synthesized through reliable synthetic routes within a short period, making them readily commercially available. These libraries encompass billions of molecules and continue to expand due to advances in synthetic accessibility (Llanos et al. 2019). While screening larger compound libraries enhances the likelihood of identifying hit compounds, it also extends screening times, thereby necessitating the scaling up of virtual screening methods (Sadybekov et al. 2022). However, scaling up 3D pharmacophore screening to accommodate billions of molecules presents significant challenges (Warr et al. 2022). Although various filtering techniques have been developed (Seidel et al. 2010), molecules that pass these methods must still undergo alignment algorithms, which ultimately determine the speed of the process. Despite substantial efforts to optimize these algorithms (Wolber et al. 2008; Permann, Seidel, and Langer 2021), the overall screening procedure remains time-intensive.
In this work, we propose using self-supervised learning to create meaningful 3D pharmacophore representations for efficient virtual screening. Our PharmacoMatch model employs a graph neural network (GNN) encoder, trained with a contrastive learning objective, to map 3D pharmacophores into an order embedding space (Ying et al. 2020), thereby enabling pharmacophore matching through vector comparisons. The embedding vectors for the screening database are computed once and then used to quickly generate a hitlist based on the query embedding. An overview of the workflow is presented in Figure 1.
Our key contributions are:
  • We develop a GNN encoder model that generates meaningful vector representations from 3D pharmacophores. The model is trained in a self-supervised manner on unlabeled data, employing a contrastive loss objective to capture the relationships between queries and targets based on their partial ordering in the learned embedding space.
  • We use the learned representation for fast virtual screening in the embedding space and evaluate the performance of our method through experiments on virtual screening benchmark datasets.
@@ 2 Related work
Alignment algorithms compute a rigid-body transformation, the pharmacophore alignment, to match a query’s pharmacophoric pattern to database ligands. A scoring function then evaluates the pharmacophore matching by considering both the number of matched features and their spatial proximity. The alignment is typically preceded by fast filtering methods that prune the search space based on pharmacophoric types, pharmacophoric point counts, and quick distance checks. Only molecules that pass these filters undergo the final, computationally expensive 3D alignment step, which is usually performed by minimizing the root mean square deviation (RMSD) between pairs of pharmacophoric points (Seidel et al. 2010; Dixon et al. 2006). The algorithm by Wolber, Dornhofer, and Langer (2006) creates smoothed histograms from the neighborhoods of pharmacophoric points for pair assignment using the Hungarian algorithm, followed by alignment with Kabsch’s method (Kabsch 1976). A recent implementation by Permann, Seidel, and Langer (2021) improves on runtime and accuracy by using a search strategy that maximizes pairs of matching pharmacophoric points. Alternatively, shape-matching algorithms like ROCS (Hawkins, Skillman, and Nicholls 2007) and Pharao (Taminau, Thijs, and De Winter 2008) model pharmacophoric points with Gaussian volumes, optimizing for volume overlap.
@@
A common approach to using machine learning for virtual screening is to train models on measured bioactivity values. However, these models are constrained by the scarcity of experimental data, which is both costly and challenging to obtain (Li et al. 2021). Unsupervised training of target-agnostic models for virtual screening avoids dependence on labeled data, but remains relatively unexplored. DrugClip (Gao et al. 2023) approaches virtual screening as a similarity matching problem between protein pockets and molecules, using a multi-modal learning approach where a protein and a molecule encoder create a shared embedding space for virtual screening. Sellner, Mahmoud, and Lill (2023) used the Schrödinger pharmacophore shape-screening score to train a transformer model on pharmacophore similarity, which is a different objective than pharmacophore matching. PharmacoNet (Seo and Kim 2023) uses instance segmentation for pharmacophore generation in protein binding sites and a graph-matching algorithm for binding pose estimation, employing deep learning for pharmacophore modeling, but not for the alignment nor matching.
Contrastive representation learning
A common approach for the extraction of vector embeddings is the use of contrastive learning frameworks. These frameworks make use of a Siamese network architecture and a contrastive loss function, where an embedding space is learned by comparing positive and negative examples (Bengio, Courville, and Vincent 2013). In the last years, the computer vision community reported great improvements in the use of self-supervised learning (SSL) frameworks, which can be seen as a special case of contrastive learning. Instead of labels, these frameworks use augmentations to create positive and negative examples during training, which allows to train models on large datasets of unlabeled data. SSL is often used for model pretraining, followed by fine-tuning through supervised learning, which is especially useful when data is limited; however, the learned representations can also be utilized without fine-tuning if no labeled data is available (Balestriero et al. 2023).
@@ 3 Preliminaries
In this work, we treat 3D pharmacophores as attributed point clouds (Mahé et al. 2006; Kriege and Mutzel 2012). A pharmacophore $P$ can be represented by a set of pharmacophoric points $P=\{(\mathbf{r}_{i},l_{i})\in\mathbb{R}^{3}\times\mathcal{L}\}_{i}$ with the Cartesian coordinates $\mathbf{r}_{i}$ and the label $l_{i}$ of the pharmacophoric point $p_{i}$. The label set $\mathcal{L}$ contains the following pharmacophoric descriptors: hydrogen bond donors (HBD) and acceptors (HBA), halogen bond donors (XBD), positive (PI) and negative electrostatic interaction sites (NI), hydrophobic interaction sites (H), and aromatic moieties (AR). Directed descriptors like HBD and HBA can be associated with a vector component, but for simplicity, we will omit this information in our study.
The pharmacophore $P$ can be represented as a complete graph $G(P)=(V_{P},E_{P},\lambda_{P})$, where $V_{P}=\{v_{1},...,v_{|P|}\}$ denotes the set of nodes with node attributes $\lambda_{P}(v_{i})=l_{i}$, and $E_{P}=V_{P}\times V_{P}$ denotes the set of edges, where $\lambda_{P}$ represents a labelling function $\lambda:V\cup E\rightarrow\mathcal{L}$ that assigns a label to the corresponding vertex $v$ or edge $e$. The edges are undirected, edge $e_{uv}$ can be identified with edge $e_{vu}$, and the label of $e_{uv}$ is the pair-wise Euclidean distance $\lambda_{P}(e_{uv})=\lVert\mathbf{r}_{u}-\mathbf{r}_{v}\rVert_{2}$ between the positions of nodes $u$ and $v$. This representation is invariant to translation and rotation.
@@
Two graphs $G_{1}=(V_{1},E_{1},\lambda_{1})$ and $G_{2}=(V_{2},E_{2},\lambda_{2})$ are isomorphic, denoted by $G_{1}\simeq G_{2}$, if there exists an edge-preserving bijection $f:V_{1}\rightarrow V_{2}$ such that $\forall(u,v)\in E_{1}:(f(u),f(v))\in E_{2}$. Additionally, we require the preservation of node and edge labels, such that $\forall v\in V_{1}:\lambda_{1}(v)=\lambda_{2}(f(v))$, and $\forall(u,v)\in E_{1}:\lambda_{1}((u,v))=\lambda_{2}((f(u),f(v)))$. Let $G_{Q}=(V_{Q},E_{Q},\lambda_{Q})$ be a query graph, $G_{T}=(V_{T},E_{T},\lambda_{T})$ a larger target graph, and $G_{H}=(V_{H},E_{H},\lambda_{H})$ a subgraph of $G_{T}$ such that $V_{H}\subseteq V_{T}$, and $E_{H}\subseteq E_{T}$. The objective of subgraph matching is to decide, whether $G_{Q}$ is subgraph isomorphic to $G_{T}$, denoted by $G_{Q}\lesssim G_{T}$, which requires the existence of a non-empty set of subgraphs =$\mathcal{H}=\{G_{H}\mid G_{H}\simeq G_{Q}\}$ that are isomorphic to $G_{Q}$.
@@
In its most general setting, pharmacophore matching seeks to match all pharmacophoric points of a query pharmacophore $P_{Q}$ with the corresponding pharmacophoric points of a larger target pharmacophore $P_{T}$.
Let $P_{H}\subseteq P_{T}$ denote a subset of the pharmacophoric points of $P_{T}$. Then $P_{Q}$ matches $P_{T}$ after alignment if there exists a bijection $g:P_{Q}\rightarrow P_{H}$ such that $\forall i\in P_{Q}:l_{i}=l_{g(i)}$ and $\lVert\mathbf{r}_{i}-\mathbf{r}_{g(i)}\rVert_{2}<r_{T}$, where $r_{T}$ is the radius of a tolerance sphere. It is thereby sufficient that query pharmacophoric points are mapped into the tolerance sphere of their target counterpart. For simplicity, we assume the same tolerance radii among all pharmacophoric points. The ultimate goal of pharmacophore matching is to retrieve molecules from a database. A matching pharmacophore is always linked to a corresponding ligand molecule via a look-up table.
@@
When represented as graphs $G_{Q}=G(P_{Q})$, $G_{H}=G(P_{H})$, and $G_{T}=G(P_{T})$, this task boils down to the node-induced subgraph matching of a query pharmacophore graph $G_{Q}$ to a target pharmacophore graph $G_{T}$. The tolerance sphere, however, weakens the requirement on edge label matching. An approximate matching $\lambda_{Q}((u,v))\approx\lambda_{H}((f(u),f(v)))$ is sufficient if the difference between $\lambda_{Q}((u,v))$ and $\lambda_{H}((f(v),f(u)))$ is less than $2r_{T}$, where $r_{T}$ represents the tolerance radius of each pharmacophoric point. This ensures that the query points fall within the tolerance spheres of the target points (compare Figure 2). Our problem formulation of pharmacophore matching relies on relative distances instead of the absolute positioning of pharmacophoric features and is therefore independent of prior alignment.
@@ 4 Methodology
Figure 3: (a.) The embedding model learns an order embedding space by comparing augmented pharmacophores. (b) Illustration of the embedding space, where pharmacophores matching a query are positioned to the upper right.
@@
In the following we introduce PharmacoMatch, a novel contrastive learning framework with the aim to encode query-target relationships of 3D pharmacophores into an embedding space. We propose to train a GNN encoder model in a self-supervised fashion, as illustrated in Figure 3. Our model is trained on approximately 1.2 million unlabeled small molecules from the ChEMBL database (Davies et al. 2015; Zdrazil et al. 2023) and learns pharmacophore matching solely from augmented examples, comparing positive and negative pairs of query and target pharmacophore graphs, while optimizing an order embedding loss to extract relevant matching patterns.
@@
To span the pharmaceutical compound space, we download a set of drug-like molecules sourced from the ChEMBL (2024) website in the form of Simplified Molecular Input Line Entry System (SMILES) strings (Weininger 1988) and curate an unlabeled dataset using the open-source Chemical Data Processing Toolkit (CDPKit) (Seidel 2024) (see Appendix A.1 for details). After an initial data clean-up, which includes the removal of solvents and counter ions, adjustment of protonation states to a physiological pH, and elimination of duplicate structures, the dataset contains approximately 1.2 million small molecules. To ensure a zero-shot setting in our validation experiments, we remove all molecules from the training data that also appear in the test sets. Finally, we generate a low-energy 3D conformation and the corresponding pharmacophore for each ligand.
@@
We represent the node labels $\lambda_{P}(V_{P})$ of a given pharmacophore graph $G(P)=(V_{P},E_{P},\lambda_{P})$ as one-hot-encoded (OHE) feature vectors $\mathbf{h}=(\mathbf{h}_{1},...,\mathbf{h}_{|P|})$. We employ a distance encoding to represent pair-wise distances, which was inspired by the SchNet architecture (Schütt et al. 2018). The edge attributes of edge $e_{uv}$ are derived from the edge label $\lambda_{P}(e_{uv})$ and represented by a radial basis function $\mathbf{e}_{k}(\mathbf{r}_{u}-\mathbf{r}_{v})=\exp(-\beta(\lVert\mathbf{r}_{u}-\mathbf{r}_{v}\rVert_{2}-{\mu}_{k})^{2})$, where centers $\mu_{k}$ were taken from a uniform grid of $K$ points between zero and the distance cutoff at 10 Å, and the smoothing factor $\beta$ represents a hyperparameter. To this end, the pharmacophore $P$ is represented by a data point $\mathbf{x}=[\mathbf{h},\mathbf{e}]$ which is a tuple of the feature matrix $\mathbf{h}\in\mathbb{R}^{|P|\times L}$ and the distance-encodings $\mathbf{e}\in\mathbb{R}^{(|P|\times|P|)\times K}$.
GNN encoder architecture
The encoder input is the pharmacophore graph representation $\mathbf{x}=[\mathbf{h},\mathbf{e}]$, with the feature matrix $\mathbf{h}$ and the edge attributes $\mathbf{e}$. Node feature embeddings are generated by initially passing the OHE feature matrix through a single dense layer without an activation function. We then update the node representations through message passing using the edge-conditioned convolution operator (NNConv) by Gilmer et al. (2017); Simonovsky and Komodakis (2017), which was originally designed for representation learning on point clouds and 3D molecules, to aggregate distance information into the learned node representations (see Appendix A.3 for details). We find that DenseNet-style skip-connections (Huang et al. 2017) are beneficial for learning robust representations. Graph-level read-out is achieved by additive pooling of the updated feature matrix $\mathbf{h}\in\mathbb{R}^{|P|\times m}$ into a graph representation $\mathbf{q}\in\mathbb{R}^{m}$, which is then projected to the final output embedding $\mathbf{z}\in\mathbb{R}^{D}_{+}$ by a multi-layer perceptron. The employed loss function requires to map the final representation to the non-negative real number space. We accomplish this by using the absolute values of the learnable weights for the last linear transformation immediately after the final ReLU unit (see Appendix A.4 for details).
@@
In order to encode query-targets relationship of pharmacophores into the embedding space, we employ the loss function by Ying et al. (2020). The key insight is that subgraph relationships can be effectively encoded in the geometry of an order embedding space through a partial ordering of the corresponding vector embeddings. Let $\mathbf{z}_{Q}$ the embedding of graph $G_{Q}$, $\mathbf{z}_{T}$ the embedding of graph $G_{T}$, and $f_{\Theta}:\mathcal{G}\rightarrow\mathbb{R}^{D}_{+}$ a GNN encoder to map pharmacophore graphs $\mathcal{G}$ to embedding vectors $\mathbf{z}\in\mathbb{R}^{D}_{+}$. The partial ordering $\mathbf{z}_{Q}\preceq\mathbf{z}_{T}$ reflects, whether $G_{Q}$ is subgraph isomorphic to $G_{T}$:
$\mathbf{z}_{Q}[i]\leq\mathbf{z}_{T}[i],\;\forall i\in\{1,...,D\}\;\mathrm{iff}\;G_{Q}\lesssim G_{T}$ (1)
The following max-margin objective can be used to train the GNN encoder $f_{\Theta}$ on this relation:
@@
The penalty function $E:\mathbb{R}^{D}_{+}\times\mathbb{R}^{D}_{+}\rightarrow\mathbb{R_{+}}$ reflects violation of the partial ordering on the embedding vector pair:
$E(\mathbf{z}_{Q},\mathbf{z}_{T})=\lVert\max\{\mathbf{0},\mathbf{z}_{Q}-\mathbf{z}_{T}\}\rVert_{2}^{2}$ (3)
$Pos$ is the set of positive pairs per batch, these are pairs of query $\mathbf{z}_{Q}$ and target graph embedding $\mathbf{z}_{T}$ with a subgraph-supergraph relationship, and $Neg$ is the set of negative examples, these are pairs of query and target embedding vectors that violate this relationship. The positive and negative pairs are generated on-the-fly via augmentation during training.
Figure 4: Augmentation strategies for model training involve generating positive and negative query-target pairs on-the-fly by combining node deletion with varying degrees of node displacement. Negative pairs are also created by shuffling the batch, mapping query pharmacophores to random target pharmacophores.

### v2

# PharmacoMatch: Efficient 3D Pharmacophore Screening via Neural Subgraph Matching

@@ Abstract
The increasing size of screening libraries poses a significant challenge for the development of virtual screening methods for drug discovery, necessitating a re-evaluation of traditional approaches in the era of big data. Although 3D pharmacophore screening remains a prevalent technique, its application to very large datasets is limited by the computational cost associated with matching query pharmacophores to database molecules. In this study, we introduce PharmacoMatch, a novel contrastive learning approach based on neural subgraph matching. Our method reinterprets pharmacophore screening as an approximate subgraph matching problem and enables efficient querying of conformational databases by encoding query-target relationships in the embedding space. We conduct comprehensive investigations of the learned representations and evaluate PharmacoMatch as pre-screening tool in a zero-shot setting. We demonstrate significantly shorter runtimes and comparable performance metrics to existing solutions, providing a promising speed-up for screening very large datasets.
@@ 1 Introduction
The immense scale of the chemical space, covering over $10^{60}$ small organic molecules (Virshup et al., 2013), makes the identification of molecules that bind to a protein target particularly challenging. Virtual screening methods are important tools in computer-aided drug discovery, addressing this complexity and supporting medicinal chemists navigate large molecular databases in search of potential hit compounds (Sliwoski et al., 2014). Among these methods, an established approach is 3D pharmacophore screening. Pharmacophores represent an ensemble of steric and electronic molecular features that is necessary to ensure an optimal interaction with a specific biological target (Wermuth et al., 1998). A pharmacophore query can, for instance, be generated from the interaction profile of a ligand-protein complex and then aligned positionally with the three-dimensional conformations of compounds in a database (Wolber & Langer, 2005). These molecules are subsequently ranked based on their agreement with the pharmacophore query, with those exhibiting similar pharmacophoric patterns retrieved as potential hit compounds (Wolber et al., 2006).
@@
An important development of the last years has been the emergence of make-on-demand libraries, such as Enamine REAL (Shivanyuk et al., 2007). These libraries contain billions of commercially available compounds that can be rapidly synthesized and are continuously expanding, driven by advances in synthetic accessibility (Llanos et al., 2019). Screening larger libraries increases the chances of identifying hits, but it also brings the trade-off of longer screening times. Scaling up 3D pharmacophore screening to handle billions of molecules is challenging due to the computational cost of pharmacophore alignment (Warr et al., 2022). Despite efforts to optimize these algorithms (Wolber et al., 2008; Permann et al., 2021) and the development of various pre-filtering techniques (Seidel et al., 2010), the alignment step remains a critical bottleneck.
In this work, we propose to perform 3D pharmacophore screening by using learned representations. Specifically, our PharmacoMatch model employs a graph neural network (GNN) encoder to map 3D pharmacophores into an order embedding space (Ying et al., 2020), and predicts pharmacophore matching through vector comparisons. The embedding vectors for the screening database are computed once and then used to quickly generate a hitlist based on the query embedding (Figure 1). Our key contributions are as follows.
  • We conceptualize pharmacophore matching as a representation learning problem, enabling the use of order embeddings for efficient and scalable pre-screening.
  • We develop a GNN encoder that generates meaningful vector representations from 3D pharmacophores. The model is trained in a self-supervised manner on unlabeled data, employing a contrastive loss objective to capture partial ordering relationships between queries and targets in the learned embedding space. We design augmentation strategies specifically suited for the task of pharmacophore matching.
  • We thoroughly analyze the learned embeddings and validate the practical utility of our approach as an effective pre-screening tool through experiments on virtual screening benchmark datasets.
We empirically demonstrate that our method is significantly faster than existing solutions, offering a practical advantage for screening billion-compound libraries. A promising direction for further improving efficiency is combining our approach with approximate retrieval techniques. Our work thus represents a key step towards vector databases for virtual screening. Additionally, our reformulation of pharmacophore screening as neural subgraph matching directly captures its subgraph matching nature. To the best of our knowledge, we are the first to introduce order embeddings in virtual screening, providing a novel perspective for the field. Furthermore, our use of self-supervised learning presents a promising strategy for addressing data scarcity in drug discovery. We hope that the combined insights from our study contribute to advancing representation learning for virtual screening applications.
@@ 2 Related work
Alignment algorithms compute a rigid-body transformation, the pharmacophore alignment, to match a query’s pharmacophoric feature pattern to database ligands. A scoring function then evaluates the pharmacophore matching by considering both the number of matched features and their spatial proximity. The alignment is typically preceded by fast filtering methods that prune the search space based on feature types, feature point counts, and quick distance checks. Only molecules that pass these filters undergo the final, computationally expensive 3D alignment step, which is usually performed by minimizing the root mean square deviation (RMSD) between pairs of pharmacophoric feature points (Dixon et al., 2006; Seidel et al., 2010). The algorithm by Wolber et al. (2006) creates smoothed histograms from the neighborhoods of pharmacophoric feature points for pair assignment using the Hungarian algorithm, followed by alignment with Kabsch’s method (Kabsch, 1976). A recent implementation by Permann et al. (2021) improves on runtime and accuracy by using a search strategy that maximizes pairs of matching pharmacophoric feature points. Alternatively, shape-matching algorithms like ROCS (Hawkins et al., 2007) and Pharao (Taminau et al., 2008) model chemical features by Gaussian volumes, optimizing for volume overlap.
@@
A common approach to using machine learning for virtual screening is to train models on measured bioactivity values. However, these models are constrained by the scarcity of experimental data, which is both costly and challenging to obtain (Li et al., 2021). Unsupervised training of target-agnostic models for virtual screening avoids dependence on labeled data, but remains relatively unexplored. DrugClip (Gao et al., 2023), which is not based on pharmacophores, approaches virtual screening as a similarity matching problem between protein pockets and molecules, using a multi-modal learning approach where a protein and a molecule encoder create a shared embedding space for virtual screening. Sellner et al. (2023) used the Schrödinger pharmacophore shape-screening score to train a transformer model on pharmacophore similarity, which is a different objective than pharmacophore matching. PharmacoNet (Seo & Kim, 2023) is a pre-screening tool that uses instance segmentation for pharmacophore generation in protein binding sites and a graph-matching algorithm for binding pose estimation. They employ deep learning for pharmacophore modeling, but not for the alignment nor matching.
@@ 3 Preliminaries
In this work, we treat 3D pharmacophores as attributed point clouds (Mahé et al., 2006; Kriege & Mutzel, 2012). A pharmacophore $P$ can be represented by a set of pharmacophoric feature points $P=\{(\mathbf{r}_{i},d_{i})\in\mathbb{R}^{3}\times\mathcal{D}\}_{i}$ with the Cartesian coordinates $\mathbf{r}_{i}$ and the descriptor $d_{i}$ of the pharmacophoric feature point $p_{i}$. The descriptor set $\mathcal{D}$ covers the following pharmacophoric feature types: hydrogen bond donors (HBD) and acceptors (HBA), halogen bond donors (XBD), positive (PI) and negative electrostatic interaction sites (NI), hydrophobic interaction sites (H), and aromatic moieties (AR). Directed feature types like HBD and HBA can be associated with a vector component, but for simplicity, we will omit this information in our study. We further denote the set of pair-wise distances between feature points as =$\mathcal{R}=\{\|\mathbf{r}_{i}-\mathbf{r}_{j}\|_{2}\mid 1\leq i,j\leq|P|\}$. The pharmacophore $P$ can be represented as a complete graph $G(P)=(V_{P},E_{P},\lambda_{P})$, where $V_{P}=\{v_{1},...,v_{|P|}\}$ denotes the set of nodes with node attributes $\lambda_{P}(v_{i})=l_{i}$, and $E_{P}=V_{P}\times V_{P}$ denotes the set of edges, with the edge attribute of $e_{ij}$ defined as the pair-wise Euclidean distance $\lambda_{P}(e_{ij})=\lVert\mathbf{r}_{i}-\mathbf{r}_{j}\rVert_{2}$ between the positions of nodes $i$ and $j$. The edges are undirected, edge $e_{ij}$ can be identified with edge $e_{ji}$. The label set $\mathcal{L}=\mathcal{D}\cup\mathcal{R}$ is the union of the feature descriptor set and the set of pair-wise distances, and $\lambda_{P}$ represents a labelling function $\lambda:V\cup E\rightarrow\mathcal{L}$ that assigns a label to the corresponding vertex $v$ or edge $e$. This representation is invariant to translation and rotation.
@@
Two graphs $G_{1}=(V_{1},E_{1},\lambda_{1})$ and $G_{2}=(V_{2},E_{2},\lambda_{2})$ are isomorphic, denoted by $G_{1}\simeq G_{2}$, if there exists an edge-preserving bijection $f:V_{1}\rightarrow V_{2}$ such that $\forall(u,v)\in E_{1}:(f(u),f(v))\in E_{2}$. Additionally, we require the preservation of node and edge labels, such that $\forall v\in V_{1}:\lambda_{1}(v)=\lambda_{2}(f(v))$, and $\forall(u,v)\in E_{1}:\lambda_{1}((u,v))=\lambda_{2}((f(u),f(v)))$. Let $G_{Q}=(V_{Q},E_{Q},\lambda_{Q})$ be a query graph, $G_{T}=(V_{T},E_{T},\lambda_{T})$ a larger target graph, and $G_{H}=(V_{H},E_{H},\lambda_{H})$ a subgraph of $G_{T}$ such that $V_{H}\subseteq V_{T}$, and $E_{H}\subseteq E_{T}$. The objective of subgraph matching is to decide, whether $G_{Q}$ is subgraph isomorphic to $G_{T}$, denoted by $G_{Q}\lesssim G_{T}$, requiring the existence of a non-empty set of subgraphs =$\mathcal{H}=\{G_{H}\mid G_{H}\simeq G_{Q}\}$ that are isomorphic to $G_{Q}$.
@@
In its most general setting, pharmacophore matching seeks to match all pharmacophoric feature points of a query pharmacophore $P_{Q}$ with the corresponding feature points of a larger target pharmacophore $P_{T}$. Let $P_{H}\subseteq P_{T}$ denote a subset of the feature points of $P_{T}$. Then $P_{Q}$ matches $P_{T}$ after alignment if there exists a bijection $g:P_{Q}\rightarrow P_{H}$ such that $\forall i\in P_{Q}:d_{i}=d_{g(i)}$ and $\lVert\mathbf{r}_{i}-\mathbf{r}_{g(i)}\rVert_{2}<r_{T}$, where $r_{T}$ is the radius of a tolerance sphere. It is thereby sufficient that query feature points are mapped into the tolerance sphere of their target counterpart. For simplicity, we assume the same tolerance radii among all pharmacophoric feature points. The ultimate goal of pharmacophore matching is to retrieve molecules from a database. A matching pharmacophore is always linked to a corresponding ligand molecule via a look-up table.
@@
When represented as graphs $G_{Q}=G(P_{Q})$, $G_{H}=G(P_{H})$, and $G_{T}=G(P_{T})$, this task boils down to the node-induced subgraph matching of a query pharmacophore graph $G_{Q}$ to a target pharmacophore graph $G_{T}$. The tolerance sphere, however, weakens the requirement on edge label matching. An approximate matching $\lambda_{Q}((u,v))\approx\lambda_{H}((f(u),f(v)))$ is sufficient if the difference between $\lambda_{Q}((u,v))$ and $\lambda_{H}((f(v),f(u)))$ is less than $2r_{T}$, where $r_{T}$ represents the tolerance radius of each pharmacophoric feature point. This ensures that the query points fall within the tolerance spheres of the target points (compare Figure 2). Our problem formulation of pharmacophore matching relies on relative distances instead of the absolute positioning of pharmacophoric features and is therefore independent of prior alignment.
@@ 4 Methodology
In the following we introduce PharmacoMatch, a novel contrastive learning framework with the aim to encode query-target relationships of 3D pharmacophores into an embedding space. We propose to train a GNN encoder model in a self-supervised fashion, as illustrated in Figure 3. Our model is trained on approximately 1.2 million unlabeled small molecules from the ChEMBL database (Davies et al., 2015; Zdrazil et al., 2023) and learns pharmacophore matching solely from augmented examples, comparing positive and negative pairs of query and target pharmacophore graphs, while optimizing an order embedding loss to extract relevant matching patterns.
@@
To span the pharmaceutical compound space, we download a set of drug-like molecules sourced from the ChEMBL (2024) website in the form of Simplified Molecular Input Line Entry System (SMILES) strings (Weininger, 1988) and curate an unlabeled dataset using the open-source Chemical Data Processing Toolkit (CDPKit) (Seidel, 2024) (see Appendix A.1 for details). After an initial data clean-up, which includes the removal of solvents and counter ions, adjustment of protonation states to a physiological pH, and elimination of duplicate structures, the dataset contains approximately 1.2 million small molecules. To ensure a zero-shot setting in our validation experiments, we remove all molecules from the training data that also appear in the test sets. Finally, we generate a low-energy 3D conformation and the corresponding pharmacophore for each ligand.
@@
We represent the node labels $\{\lambda_{P}(v_{1}),...,\lambda_{P}(v_{|P|})\}$ of a given pharmacophore graph $G(P)=(V_{P},E_{P},\lambda_{P})$ as one-hot-encoded (OHE) feature vectors $\mathbf{h}=(\mathbf{h}_{1},...,\mathbf{h}_{|P|})$. We employ a distance encoding to represent pair-wise distances, which was inspired by the SchNet architecture (Schütt et al., 2018). The edge attributes of edge $e_{uv}$ are derived from the edge label $\lambda_{P}(e_{uv})$ and represented by a radial basis function $\mathbf{e}_{k}(\mathbf{r}_{u}-\mathbf{r}_{v})=\exp(-\beta(\lVert\mathbf{r}_{u}-\mathbf{r}_{v}\rVert_{2}-{\mu}_{k})^{2})$, where centers $\mu_{k}$ were taken from a uniform grid of $K$ points between zero and the distance cutoff at 10 Å, and the smoothing factor $\beta$ represents a hyperparameter. To this end, the pharmacophore $P$ is represented by a data point $\mathbf{x}=[\mathbf{h},\mathbf{e}]$ which is a tuple of the feature matrix $\mathbf{h}\in\mathbb{R}^{|P|\times|\mathcal{D}|}$ and the distance-encodings $\mathbf{e}\in\mathbb{R}^{(|P|\times|P|)\times K}$.
GNN encoder
The encoder input is the pharmacophore graph representation $\mathbf{x}=[\mathbf{h},\mathbf{e}]$, with the feature matrix $\mathbf{h}$ and the edge attributes $\mathbf{e}$. Node feature embeddings are generated by initially passing the OHE feature matrix through a single dense layer without an activation function. We then update the node representations through message passing using the edge-conditioned convolution operator (NNConv) by Gilmer et al. (2017); Simonovsky & Komodakis (2017), which was originally designed for representation learning on point clouds and 3D molecules, to aggregate distance information into the learned node representations (see Appendix A.3 for details). We connect successive convolutional layers using DenseNet-style skip connections (Huang et al., 2017). Graph-level read-out is achieved by additive pooling of the updated feature matrix $\mathbf{h}\in\mathbb{R}^{|P|\times m}$ into a graph representation $\mathbf{q}\in\mathbb{R}^{m}$, which is then projected to the final output embedding $\mathbf{z}\in\mathbb{R}^{D}_{+}$ by a multi-layer perceptron. The employed loss function requires to map the final representation to the non-negative real number space. We accomplish this by using the absolute values of the learnable weights for the last linear transformation immediately after the final ReLU unit (see Appendix A.4 for details).
Figure 3: (a) The encoder model learns an order embedding space by comparing augmented pharmacophores. (b) Illustration of the embedding space, where pharmacophores matching a query are positioned to the upper right. (c) Augmentation strategies for model training involve generating positive and negative query-target pairs on-the-fly by combining node deletion with varying degrees of node displacement. Negative pairs are also created by shuffling the batch, mapping query pharmacophores to random target pharmacophores.
@@
In order to encode query-target relationships of pharmacophores into the embedding space, we employ the loss function by Ying et al. (2020). The key insight is that subgraph relationships can be effectively encoded in the geometry of an order embedding space through a partial ordering of the corresponding vector embeddings. Let $\mathbf{z}_{Q}$ the embedding of graph $G_{Q}$, $\mathbf{z}_{T}$ the embedding of graph $G_{T}$, and $f_{\Theta}:\mathcal{G}\rightarrow\mathbb{R}^{D}_{+}$ a GNN encoder to map pharmacophore graphs $\mathcal{G}$ to embedding vectors $\mathbf{z}\in\mathbb{R}^{D}_{+}$. The partial ordering $\mathbf{z}_{Q}\preceq\mathbf{z}_{T}$ reflects, whether $G_{Q}$ is subgraph isomorphic to $G_{T}$: $\mathbf{z}_{Q}[i]\leq\mathbf{z}_{T}[i],\;\forall i\in\{1,...,D\}\;\mathrm{iff}\;G_{Q}\lesssim G_{T}$. The following max-margin objective can be used to train the GNN encoder $f_{\Theta}$ on this relation:
@@
The penalty function $E:\mathbb{R}^{D}_{+}\times\mathbb{R}^{D}_{+}\rightarrow\mathbb{R_{+}}$ reflects violation of the partial ordering on the embedding vector pair: $E(\mathbf{z}_{Q},\mathbf{z}_{T})=\lVert\max\{\mathbf{0},\mathbf{z}_{Q}-\mathbf{z}_{T}\}\rVert_{2}^{2}$. $Pos$ is the set of positive pairs per batch, these are pairs of query $\mathbf{z}_{Q}$ and target graph embedding $\mathbf{z}_{T}$ with a subgraph-supergraph relationship, and $Neg$ is the set of negative examples, these are pairs of query and target embedding vectors that violate this relationship. The positive and negative pairs are generated on-the-fly via augmentation during training.